OpenStudy (anonymous):

Rhett decides to build a square room for his movie and music collection. If the area of the room is 9x2 − 6x + 1 square feet, what is the length of one side of the room? (3x + 1) feet (9x − 1) feet (3x − 1) feet (9x + 1) feet

OpenStudy (nurali):

9x² - 6x = - 1 x² - 1/3x = - 1/9 + (- 1/3)² x² - 1/3x = - 1/9 + 1/3 x - 1/3 = 0 = x - 1/3, = 3x - 1

OpenStudy (goformit100):

is it 9^2 or 9x2

OpenStudy (anonymous):

9x^2

OpenStudy (anonymous):

of course @goformit100 it is 9x^2